1. Calculate the molar solubility of Ag ion in AgCl? (Ksp = 1.6x10-10)

AgCl (s) ↔ Ag+ (aq) + Cl- (aq)

2.Ka for acetic acid is 1.75x10-5. Find Kb for acetate ion at 25oC. (Kw= 1.0 x 10-14)


1
Expert's answer
2021-03-05T03:30:33-0500

1.We know that AgCl is aqueous solution dissociate as and (Ksp = 1.6x10-10

AgCl (s) ↔ Ag+ (aq) + Cl- (aq)

take molar solubility as S and it dissociate into two ions as Ag+ and Cl-

"K_{sp}= S\times S"

"S^2= K_{sp}"

"S^2=1.6 \times 10^{-10}"

"S=1.26\times10^{-5}"


2 In the question it is given that

Ka for acetic acid is 1.75x10-5 and also Kw= 1.0 x 10-14

we know that

"K_a \times K_b= K_w"

"1.75\times 10^{-5} \times K_b= 1.0 \times 10^{-14}"

"K_b=5.7 \times 10^{-8}"